OpenStudy (anonymous):

I have an answer I just need someone to compare with Evaluate 15-3(2+(3-6)/8-4)-3^2

OpenStudy (anonymous):

\[15-3(2+\frac{ 3-6 }{8-4 })-3^{2}\]

OpenStudy (whpalmer4):

What do you get for an answer?

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

I got 6

OpenStudy (whpalmer4):

\[15-3(2+\frac{(3-6)}{(8-4)}) - 3^2 = 15-6 -3(\frac{-3}{4}) -9 = \frac{9}{4}\]

OpenStudy (anonymous):

I see how you got that can you just see why mine is not right\[15-3(2+\frac{ 3-6 }{ 8-4) }-3^{2}=12(2+(\frac{ 3-6 }{ 8-4})-3^{2}=12(1.25)-3^{2}\] Answer 6

OpenStudy (whpalmer4):

Because it isn't (15-3)(stuff), it's 15 - 3(stuff)

OpenStudy (anonymous):

oh ok

OpenStudy (whpalmer4):

You could write it as \[-3(2+\frac{3-6}{8-4}) + 15 - 9\]and then you wouldn't be so likely to make that mistake.